17-21 Five Measures of Solvency or Profitability The balance sheet for Garcon Inc. at the end...

17-21

Five Measures of Solvency or Profitability

The balance sheet for Garcon Inc. at the end of the current fiscal year indicated the following:

Bonds payable, 10% $1,700,000
Preferred $10 stock, $100 par 196,000
Common stock, $11 par 689,920.00

Income before income tax was $391,000, and income taxes were $57,800 for the current year. Cash dividends paid on common stock during the current year totaled $125,440. The common stock was selling for $50 per share at the end of the year.

Determine each of the following. Round answers to one decimal place, except for dollar amounts which should be rounded to the nearest whole cent. Use the rounded answers for subsequent requirements, if required.

a. Times interest earned ratio times
b. Earnings per share on common stock $
c. Price-earnings ratio
d. Dividends per share of common stock $
e. Dividend yield %
Answer #1

a. Times interest earned ratio = (Income before income tax + Interest expense)/Interest expense

= ($391,000+1,700,000*10%)/1,700,000*10%

= 3.3 times

b. Earnings per share on common stock = (Net income - preferred dividend)/average common stock outstanding

= (391,000-57800-19600)/(689,920/11)

= $5 per share

c. Price-earnings ratio = market price / Earnings per share on common stock

= 50 / 5 = 10

d. Dividends per share of common stock = 125,440 / (689,920/11)

= $2 per share

e. Dividend yield = Dividends per share of common stock / market price

= 2/50 = 4%